# SegVol: Universal and Interactive Volumetric Medical Image Segmentation

### v3

@@ Acknowledgement
Table 1: Supervised fine-tuning datasets and validation datasets.
@@
Table 2: Data availability for supervised fine-tuning datasets and validation datasets.
@@
Table 3: Internal validation results of 3DUX-NET, SwinUNETR, nnU-Net and SegVol on the test set of supervised fine-tuning datasets in term of Dice score.
@@
Portal/splenic vein 0.7207 (0.6211, 0.7588) 0.6656 (0.5888, 0.6982) 0.7964 (0.7524, 0.8582) 0.7188 (0.7128, 0.7569)
@@
00footnotetext: Note: Dice scores are displayed as Median values (First quartile, Third quartile).
Table 4: Generalization experiment results of SegVol on the MRI set of CHAOS[28, 30, 29] dataset in term of Dice score.
Method Liver Spleen Left Kidney Right Kidney
00footnotetext: Note: Dice scores are displayed as Median values (First quartile, Third quartile).
Table 5: External validation results of SAM(Point), SAM(Bbox), SAM-MED2D, SAM-MED3D, MedSAM and SegVol on the external validation datasets in term of Dice score.
@@
ULS23(DeepLesion3D) 0.3686 (0.0855, 0.7680) 0.7473 (0.6817, 0.8063) 0.3258 (0.1325, 0.5707) 0.2386 (0.1045, 0.4372) 0.7680 (0.7103, 0.8160) 0.7065 (0.6247, 0.7782)
ULS23(Bone) 0.4461 (0.2349, 0.6676) 0.6671 (0.5854, 0.7443) 0.1947 (0.0898, 0.3969) 0.4447 (0.1481, 0.7026) 0.6896 (0.6128, 0.7530) 0.6920 (0.6097, 0.7702)
ULS23(Pancreas) 0.0675 (0.0471, 0.1237) 0.5579 (0.4911, 0.6111) 0.5548 (0.4862, 0.6382) 0.5526 (0.3287, 0.6786) 0.6561 (0.5857, 0.7143) 0.7265 (0.6722, 0.7776)
00footnotetext: Note: Dice scores are displayed as Median values (First quartile, Third quartile).
Algorithm 1 SegVol training loop
1:SegVol model, training image $x$, ground truth mask set $Y_{x}=\{y_{i}\}_{i=1}^{n}$, category set $S_{x}=\{s_{i}\}_{i=1}^{n}$, pseudo mask set $Z_{x}=\{z_{i}\}_{i=1}^{m}$
2:SegVol Model parameters
3:$n_{\text{pt}}\Leftarrow 6$
4:$\alpha\Leftarrow 0.1$
5:# Loop for each category of this case.
6:for $i\Leftarrow 1$ to $n$ do
7:     $f_{\text{img}}\Leftarrow$ model.ImageEncoder$(x)$
8:     $pt_{\text{spatial}}\Leftarrow$ prompt_generate$(y_{i})$
9:     $pt_{\text{semantic}}\Leftarrow$ TextTemplate$(s_{i})$
10:     $l_{\text{gt}}\Leftarrow$ 0
11:     # Loop for possible prompt composite types of ground truth mask.
12:     for $p\Leftarrow 1$ to $n_{pt}$ do
13:         # Choose one prompt composite type.
14:         $pt_{\text{spatial}}^{\prime}$, $pt_{\text{semantic}}^{\prime}\Leftarrow$ PromptStrategy($pt_{\text{spatial}}$, $pt_{\text{semantic}}$)
15:         $f_{\text{text}}\Leftarrow$ model.TextEncoder($pt_{\text{semantic}}^{\prime}$)
16:         $f_{\text{prompt}}\Leftarrow$ model.PromptEncoder($pt_{\text{spatial}}^{\prime}$, $f_{\text{text}}$)
17:         $pred_{\text{gt}}\Leftarrow$ model.Decoder($f_{img}$, $f_{\text{prompt}}$, $f_{\text{text}}$)
18:         $l_{\text{gt}}\Leftarrow l_{\text{gt}}$ + DiceLoss($pred_{\text{gt}}$, $y_{i}$) + BCELoss($pred_{\text{gt}}$, $y_{i}$)
19:     end for
20:     $l_{\text{pseudo}}\Leftarrow$ 0
21:     # Loop for several pseudo masks.
22:     for $p\Leftarrow 1$ to $n_{pt}$ do
23:         # Random select a pseudo mask of this case for training.
24:         $z_{p}\Leftarrow$ RandomSelectInt($[1,m]$)
25:         $pt_{\text{spatial}}\Leftarrow$ prompt_generate$(z_{p})$
26:         $f_{\text{prompt}}\Leftarrow$ model.PromptEncoder($pt_{\text{spatial}}$)
27:         $pred_{\text{pseudo}}\Leftarrow$ model.Decoder($f_{img}$, $f_{\text{prompt}}$)
28:         $l_{\text{pseudo}}\Leftarrow l_{\text{pseudo}}$ + DiceLoss($pred_{\text{pseudo}}$, $z_{p}$) + BCELoss($pred_{\text{pseudo}}$, $z_{p}$)
29:     end for
30:     $l\Leftarrow l_{\text{gt}}+\alpha\times l_{\text{pseudo}}$
31:     update(model, $l$)
32:end for
33:return model
Figure 1: An overview of the model architecture of the proposed SegVol. a The main structure of SegVol includes the image encoder, text encoder, prompt encoder, and mask decoder. All networks, except the text encoder, are learnable. The image encoder extracts the image embedding of volumetric input. The image embedding is fed into the decoder together with prompt embeddings to predict the segmentation mask. b Illustration of input image transformation and the prompt generation. c Zoom-out-zoom-in training: SegVol is trained on data of both global and local views. d Zoom-out-zoom-in inference: SegVol first conducts global inference and then performs local inference on the extracted ROI to refine the results.
Figure 2: Overview of the joint supervised fine-tuning dataset. The joint dataset comprises 47 important regions, with each region containing one or multiple significant anatomical structures within that spatial area. Image of human body by brgfx on Freepik[3].
Figure 3: Visualized aorta and left kidney prediction results of 3DUX-NET[33], SwinUNETR[15], nnU-Net[24] and SegVol on 4 cases from internal validation set. For the integrality of aorta and left kidney structure modeling, SegVol significantly outperforms 3DUX-NET and SwinUNETR and is comparable to nnU-Net.
Figure 4: Visualized liver and pancreas prediction results of 3DUX-NET[33], SwinUNETR[15], nnU-Net[24] and SegVol on 4 cases from internal validation set. For the modeling of pancreas, SegVol is significantly superior to other baseline methods.
Figure 5: Visualized spleen and stomach prediction results of 3DUX-NET[33], SwinUNETR[15], nnU-Net[24] and SegVol on 4 cases from internal validation set. For the consistency and stability of stomach modeling, SegVol is significantly better than other methods.
@@
Figure 6: Visualized aorta and bladder prediction results of MedSAM[38], SAM(bbox)[31], SAM-MED2D[6], SAM-MED3D[64], SAM(points)[31] and SegVol on 4 cases from external validation set.
@@
Figure 7: Visualized gall bladder and left kidney prediction results of MedSAM[38], SAM(bbox)[31], SAM-MED2D[6], SAM-MED3D[64], SAM(points)[31] and SegVol on 4 cases from external validation set.
@@
Figure 8: Visualized liver and prostate/uterus prediction results of MedSAM[38], SAM(bbox)[31], SAM-MED2D[6], SAM-MED3D[64], SAM(points)[31] and SegVol on 4 cases from external validation set.

### v5

@@ Appendix A Dataset Details and Availability
Table 4: Information of datasets involved in supervised fine-tuning and experiments.
@@
Table 5: Availability of datasets involved in supervised fine-tuning and experiments.
@@
Figure 6: Overview of the collected datasets for supervised fine-tuning. The joint dataset comprises 47 important regions, with each region containing one or multiple significant anatomical structures within that spatial area. Image of the human body by brgfx on Freepik[76].
Figure 7: The joint dataset encompasses various anatomical structures in major regions of the human body. Several volume examples are demonstrated as 2D slices and 3D shapes in the images respectively.
Appendix B Training Algorithm
Due to the complexity of the training steps, which include decoder reuse, the combination of different datasets, and cooperative training of ground-truth and pseudo labels, we abstract the core training code as Algorithm 1 and Figure 8 to clarify the training process of SegVol. As shown in Figure 8, each case (training sample) consists of an Image $\bm{x}$, a Ground Truth(GT) Mask Set $\bm{Y}$, and a Pseudo Mask Set $\bm{Z}$. The training loss of each sample consists of the ground-truth loss and the pseudo loss. The ground-truth loss is computed by inputting the image, the ground-truth mask (label), and the sampled prompt into the model, while the pseudo loss is computed by inputting the image, the pseudo label, and the pre-designed prompt into the model. Finally, the model is optimized by minimizing the weighted sum of the two losses.
Besides, we add a reinforcement branch for semantic-prompt in the mask decoder. We further compute a similarity matrix between the up-scaled embedding from the transposed convolution output and the text embedding. The element-wise multiplication of the similarity matrix with the mask prediction is applied before interpolation, after which the model outputs the masks.
Algorithm 1 SegVol training loop
1:SegVol model, training image $\bm{x}$, ground truth mask set ${\bm{Y_{x}}}=\{\bm{y}_{i}\}_{i=1}^{n}$, pseudo mask set $\bm{Z_{x}}=\{\bm{z}_{i}\}_{i=1}^{m}$
2:SegVol model parameters
3:$n\Leftarrow 6$ Number of combinations of 3 prompt types: text, point, and bbox.
4:$\alpha\Leftarrow 0.1$ Pseudo loss weight.
5:# Loop for each category of this case.
6:for $i\Leftarrow 1$ to $n$ do
7:     $\bm{f}_{\text{img}}\Leftarrow$ model.ImageEncoder$(\bm{x})$
8:     $\bm{pt}_{\text{spatial}}$, $\bm{pt}_{\text{semantic}}$, $\Leftarrow$ prompt_generate$(\bm{y}_{i})$
9:     $l_{\text{gt}}\Leftarrow$ 0
10:     # Loop for possible prompt combination types of ground truth mask.
11:     for $p\Leftarrow 1$ to $n$ do
12:         # Choose prompt combination type.
13:         $\bm{pt}_{\text{spatial}}^{\prime}$, $\bm{pt}_{\text{semantic}}^{\prime}\Leftarrow$ PromptStrategy($\bm{pt}_{\text{spatial}}$, $\bm{pt}_{\text{semantic}}$)
14:         $\bm{f}_{\text{text}}\Leftarrow$ model.TextEncoder($\bm{pt}_{\text{semantic}}^{\prime}$)
15:         $\bm{f}_{\text{prompt}}\Leftarrow$ model.PromptEncoder($\bm{pt}_{\text{spatial}}^{\prime}$, $\bm{f}_{\text{text}}$)
16:         $\bm{pred}_{\text{gt}}\Leftarrow$ model.Decoder($\bm{f}_{\text{img}}$, $\bm{f}_{\text{prompt}}$, $\bm{f}_{\text{text}}$)
17:         $l_{\text{gt}}\Leftarrow l_{\text{gt}}$ + DiceLoss($\bm{pred}_{\text{gt}}$, $\bm{y}_{i}$) + BCELoss($\bm{pred}_{\text{gt}}$, $\bm{y}_{i}$)
18:     end for
19:     $l_{\text{pseudo}}\Leftarrow$ 0
20:     # Loop for several pseudo masks.
21:     for $p\Leftarrow 1$ to $n$ do
22:         # Random select a pseudo mask of this case for training.
23:         $\bm{z}_{p}\Leftarrow$ RandomSelect($\bm{Z_{x}}$, $[1,m]$)
24:         $\bm{pt}_{\text{spatial}}\Leftarrow$ prompt_generate$(\bm{z}_{p})$
25:         $\bm{f}_{\text{prompt}}\Leftarrow$ model.PromptEncoder($\bm{pt}_{\text{spatial}}$)
26:         $\bm{pred}_{\text{pseudo}}\Leftarrow$ model.Decoder($\bm{f}_{\text{img}}$, $\bm{f}_{\text{prompt}}$)
27:         $l_{\text{pseudo}}\Leftarrow l_{\text{pseudo}}$ + DiceLoss($\bm{pred}_{\text{pseudo}}$, $\bm{z}_{p}$) + BCELoss($\bm{pred}_{\text{pseudo}}$, $\bm{z}_{p}$)
28:     end for
29:     $l\Leftarrow l_{\text{gt}}+\alpha\times l_{\text{pseudo}}$
30:     update(model, $l$)
31:end for
32:return model
Figure 8: The demonstration of the training algorithm. Specifically, each case (training sample) consists of an Image x, a Ground Truth(GT) Mask Set Y, and a Pseudo Mask Set Z. The training loss of each sample consists of the ground-truth loss and the pseudo loss. The ground-truth loss is computed by inputting the image, the ground-truth mask (label), and the sampled prompt into the model, while the pseudo loss is computed by inputting the image, the pseudo label, and the fixed prompt into the model. Finally, the model is updated by minimizing the weighted sum of the two losses.
Table 6: Complexity comparison of popular methods.
Method Total Avg. Avg. Avg.
[…]
SAM[28] 94M 1.3e+13 2.1764 0.5271
MedSAM[29] 94M 1.3e+13 2.1886 0.5371
SAM-MED2D[38] 271M 2.3e+12 3.5547 0.6668
SAM-MED3D[39] 101M 1.0e+11 0.1768 0.6200
SegVol 181M 6.7e+11 0.3283 0.8593
Appendix C Additional Experimental Analysis
Comparative experiments to compare with task-specific segmentation models.
Task-specific segmentation models mainly fall into two architectures, CNN-based models and Transformer-based models. We conduct comparative experiments with representative CNN-based models i.e. 3DUX-Net[23] and nnU-Net[22], and representative Transformer-based models i.e. SwinUNETR[20]. We conduct additional comparative experiments on the split 20% test set of the datasets. 10 segmentation tasks are selected from BTCV[52] and MSD-spleen[56] datasets, which focus on organ segmentation, and from MSD-lung, MSD-colon, and MSD-liver datasets, which focus on lesion segmentation. We train task-specific segmentation models on each dataset individually for each method.
The quantitative experimental results are summarized in Figure 9. Generally speaking, SegVol, jointly trained on 25 datasets, outperforms traditional task-specific segmentation models trained on a single dataset. Compared to these strong baselines, SegVol exhibits a narrower distribution of Dice scores across the eight tasks, indicating its robustness and good generalization ability. This mainly owes to the massive knowledge learned from diverse samples of the same categories but different datasets. SegVol depicts excellent performance on lesion tasks which are more challenging in semantic understanding and spatial locating. We present a detailed comparison to nnU-Net[22] on lesion tasks. As shown in Table 7, the average Dice score of SegVol is 14.76% higher than that of nnU-Net for lesion tasks. We visualize the prediction results of the two methods in Figure 10, which intuitively show that SegVol performs more precise segmentation of the tumors than nnU-Net. The detailed scores and visualization results are presented in Table 9 and Figure 11 12, and 13.
We analyze that there are mainly three factors that make SegVol more powerful than traditional task-specific models: 1) Massive generative pre-training on unlabeled data endows SegVol with a complete understanding of the volumetric structures and the discriminative feature representations, which is much superior to learning from a small number of samples. 2) Learning from joint datasets with semantic-prompt makes SegVol generalize better to unseen data and categories. For instance, SegVol can learn from both the ‘left kidney’ and ‘kidney’ categories based on their semantic correlation, while traditional task-specific models treat the two categories independently. 3) SegVol can be prompted with (spatial) points/bboxes, which provide a precise spatial reference, and (semantic) texts, which disambiguate the overlap of multiple categories in the same space. In contrast, traditional methods are not able to understand semantics. This ability enables SegVol to perform better than traditional methods in challenging tasks, e.g., segmenting lesions.
Figure 9: Violin plots for comparing experiment results of SegVol and task-specific methods. The vertical axis is the Dice score.
Table 7: The comparison of the average Dice score of SegVol and nnU-Net[22] across 3 lesion segmentation tasks.
Method Lung Tumor Colon Cancer Liver Tumor
nnU-Net[22] 0.5963 0.3769 0.3606
OURS 0.7122 0.6965 0.7825
Figure 10: Visualization results of SegVol and nnU-Net across 3 lesion segmentation tasks.
Table 8: Few-shot fine-tuning experiment on FLARE22[56, 57] and MSD-spleen[56]. SegVol* represents the model fine-tuned on all datasets.
Avg. Dice Score 100 epochs 200 epochs 300 epochs 400 epochs 500 epochs SegVol*
FLARE22[56, 57] 0.0463 0.4028 0.4926 0.5617 0.5567 0.8822
MSD-spleen[56] 0.7566 0.7866 0.9433 0.9454 0.9471 0.9597
@@ Appendix C Additional Experimental Analysis
Table 9: Comparative experiment results of 3DUX-NET, SwinUNETR, nnU-Net, and SegVol on the test set of supervised fine-tuning datasets in terms of Dice score. Dice scores are displayed as ‘Median values (First quartile, Third quartile)’.
@@
Portal vein & splenic vein 0.7207 (0.6211, 0.7588) 0.6656 (0.5888, 0.6982) 0.7964 (0.7524, 0.8582) 0.7188 (0.7128, 0.7569)
@@
Figure 11: Visualized aorta and left kidney prediction results of 3DUX-NET[23], SwinUNETR[20], nnU-Net[22] and SegVol on 4 cases from the split test set. For the integrality of aorta and left kidney structure modeling, SegVol significantly outperforms 3DUX-NET and SwinUNETR and is comparable to nnU-Net.
Figure 12: Visualized liver and pancreas prediction results of 3DUX-NET[23], SwinUNETR[20], nnU-Net[22] and SegVol on 4 cases from the split test set. For the modeling of pancreas, SegVol is significantly superior to other baseline methods.
Figure 13: Visualized spleen and stomach prediction results of 3DUX-NET[23], SwinUNETR[20], nnU-Net[22] and SegVol on 4 cases from the split test set. For the consistency and stability of stomach modeling, SegVol is significantly better than other methods.
Supplement results for comparative experiments on SAM-like interactive segmentation methods.
In this work, we compare SegVol with 5 SAM-like interactive segmentation methods on AMOS22[44], ULS23[74], and SegTHOR[75] datasets. The detailed records of Dice score are demonstrated in Table 10. The visualization results are shown in Figure 14, Figure 15, Figure 16. In this experiment, SegVol is driven by ‘bbox+text’ prompt. We demonstrate the consistency results among different prompt settings of SegVol in Figure 17, which is also conducted on AMOS22[44] and ULS23[74]. Relatively poor text prompt results in ULS23 are due to the unclear category of the dataset.
We also compare the Total Parameters, the average Multiply-Accumulates(MACs), and the average Time required to process a case of the different SAM-like methods, as shown in Table6. The comparison indicates that our method takes less computational cost while achieving much better performance. Note that when calculating MACs Per Case and Time Per Case, the slice-by-slice calculation of the 2D method and the scanning process of the 3D method are accumulated respectively. SAM-MED3D[39] only processes volume with a size of $128\times 128\times 128$. The experiments are implemented on the validation set of AMOS22[44], and the setting is the same as that in Sec. 3.2
Table 10: Comparative experiment results of SAM(Point), SAM(Bbox), SAM-MED2D, SAM-MED3D, MedSAM, and SegVol in terms of Dice score. Dice scores are displayed as ‘Median values (First quartile, Third quartile)’. The three columns in the table are, in order, the AMOS22[44], ULS23[74], and SegTHOR[75] datasets.
@@
DeepLesion3D 0.3686 (0.0855, 0.7680) 0.7473 (0.6817, 0.8063) 0.3258 (0.1325, 0.5707) 0.2386 (0.1045, 0.4372) 0.7680 (0.7103, 0.8160) 0.7065 (0.6247, 0.7782)
BoneLesion 0.4461 (0.2349, 0.6676) 0.6671 (0.5854, 0.7443) 0.1947 (0.0898, 0.3969) 0.4447 (0.1481, 0.7026) 0.6896 (0.6128, 0.7530) 0.6920 (0.6097, 0.7702)
PancreasLesion 0.0675 (0.0471, 0.1237) 0.5579 (0.4911, 0.6111) 0.5548 (0.4862, 0.6382) 0.5526 (0.3287, 0.6786) 0.6561 (0.5857, 0.7143) 0.7265 (0.6722, 0.7776)
Aorta 0.2744 (0.2289, 0.3488) 0.3894 (0.3356, 0.4325) 0.8077 (0.7693, 0.8289) 0.7703 (0.7090, 0.8133) 0.3278 (0.2786, 0.3650) 0.8439 (0.8246, 0.8673)
Esophagus 0.0348 (0.0250, 0.0439) 0.2046 (0.1416, 0.2455) 0.3578 (0.2545, 0.4281) 0.6394 (0.5477, 0.7267) 0.2196 (0.1633, 0.2630) 0.7201 (0.6512, 0.7598)
Heart 0.6695 (0.6180, 0.7419) 0.8876 (0.8652, 0.9073) 0.6012 (0.4050, 0.6922) 0.8326 (0.8059, 0.8536) 0.8924 (0.8782, 0.9099) 0.8172 (0.7664, 0.8609)
Trachea 0.9147 (0.6688, 0.9545) 0.1611 (0.1164, 0.2218) 0.8306 (0.7501, 0.8642) 0.8485 (0.8110, 0.8911) 0.1261 (0.1012, 0.1648) 0.8807 (0.8571, 0.9044)
@@
Figure 14: Visualized aorta and bladder prediction results of MedSAM[29], SAM(bbox)[28], SAM-MED2D[38], SAM-MED3D[39], SAM(points)[28] and SegVol on 4 cases from split test data.
@@
Figure 15: Visualized gall bladder and left kidney prediction results of MedSAM[29], SAM(bbox)[28], SAM-MED2D[38], SAM-MED3D[39], SAM(points)[28] and SegVol on 4 cases from split test data.
@@
Figure 16: Visualized liver and prostate/uterus prediction results of MedSAM[29], SAM(bbox)[28], SAM-MED2D[38], SAM-MED3D[39], SAM(points)[28] and SegVol on 4 cases from split test data.
Figure 17: The bar chart illustrates the consistency of SegVol’s performance across different prompt types on the validation set of AMOS22[44] and ULS23[74].
Supplement results for ablation studies on zoom-out-zoom-in mechanism.
We conduct ablation study on zoom-out-zoom-in mechanism on the split 20% test data of AMOS22[44] dataset. As shown in Table 11, the zoom-out-zoom-in mechanism achieves higher Dice scores compared to the resize and sliding window strategies in 15 organ categories. In the aspect of inference efficiency, the zoom-out-zoom-in mechanism is also very competitive and quite close to the simple resize method.
Table 11: Dice score and inference time results of ablation study on zoom-out-zoom-in mechanism.
Category Mechanism Dice Score Avg.$\uparrow$ Time Per Case Avg.$\downarrow$
Arota Resize 0.6375 64 ms
Zoom-out-zoom-in 0.8853 341 ms
Sliding window 0.8715 3452 ms
Bladder Resize 0.3229 65 ms
Zoom-out-zoom-in 0.6530 135 ms
Sliding window 0.5146 3098 ms
Duodenum Resize 0.2821 64 ms
Zoom-out-zoom-in 0.6414 168 ms
Sliding window 0.5013 3442 ms
Esophagus Resize 0.1778 62 ms
Zoom-out-zoom-in 0.4530 180 ms
Sliding window 0.3372 3186 ms
Gall bladder Resize 0.3558 63 ms
Zoom-out-zoom-in 0.6830 140 ms
Sliding window 0.5758 3220 ms
Left adrenal gland Resize 0.0937 64 ms
Zoom-out-zoom-in 0.4542 134 ms
Sliding window 0.2080 2935 ms
Left kidney Resize 0.6472 64 ms
Zoom-out-zoom-in 0.9362 177 ms
Sliding window 0.9046 3440 ms
Liver Resize 0.8252 62 ms
Zoom-out-zoom-in 0.9593 287 ms
Sliding window 0.9516 3450 ms
Pancreas Resize 0.3444 65 ms
Zoom-out-zoom-in 0.7292 159 ms
Sliding window 0.6206 3447 ms
Postcava Resize 0.5882 63 ms
Zoom-out-zoom-in 0.7901 296 ms
Sliding window 0.7769 3453 ms
Prostate/uterus Resize 0.4072 65 ms
Zoom-out-zoom-in 0.6380 136 ms
Sliding window 0.5500 3034 ms
Right adrenal gland Resize 0.1660 64 ms
Zoom-out-zoom-in 0.5260 141 ms
Sliding window 0.3624 3442 ms
Right kidney Resize 0.6570 64 ms
Zoom-out-zoom-in 0.8909 175 ms
Sliding window 0.9017 3441 ms
Spleen Resize 0.6827 89 ms
Zoom-out-zoom-in 0.8942 199 ms
Sliding window 0.8941 3481 ms
Stomach Resize 0.5752 63 ms
Zoom-out-zoom-in 0.8136 181 ms
Sliding window 0.8238 3446 ms
Generalization performance of SegVol on MRI.
We discuss the generalization performance of SegVol on an external MRI dataset. We collect 60 MRI scans annotated with 4 key organ categories from CHAOS[40, 41, 42] dataset and evaluate the generalization ability to unseen modality of SegVol. It achieves median Dice scores of 85.70%, 80.09%, 80.04%, and 81.46% for liver, spleen, left kidney, and right kidney, respectively. This generalization result demonstrates the robustness of SegVol in the face of completely unseen modality data. The detailed scores and visualization results are presented in Table 12 and Figure 18.
Few-shot fine-tuning experiment on small datasets.
To evaluate the few-shot learning ability of our model, we conduct the few-shot fine-tuning experiment on small datasets, FLARE22[56, 57] (40 training cases) and MSD-spleen[56] (32 training cases). Table 8 demonstrates that 1) fine-tuning SegVol on dozens of samples works well on easy datasets such as MSD-spleen, in which the few-shot learning performance is close to the joint fine-tuning on all datasets; 2) for challenging datasets such as FLARE22, fine-tuning on all datasets can achieve much better performance.
Table 12: Generalization experiment results of SegVol on the MRI set of CHAOS[40, 41, 42] dataset in term of Dice score. Dice scores are displayed as ‘Median values (First quartile, Third quartile)’.
Method Liver Spleen Left Kidney Right Kidney